liquids A and B form an ideal solution at 25 degrees Celsius . the vapor pressure of pure A is 300 torr and that of B is 125 torr . calculate the mole fraction of B in the solution when the vapor pressure of both A and B are equal at 25 degrees Celsius .
1
Expert's answer
2020-11-17T10:44:28-0500

PA=XAP0A

PB=XBP0B

Where, PA=new vapor pressure of A

PB=new vapor pressure of B

XA=mole fraction of A

XB=mole fraction of B

P0A=Vapor pressure of pure A

P0B=Vapor pressure of pure B

P0A=300torr

P0B=125torr

According to the question, P0B= PB

Therefore, XAP0A=XBP0B

If mole fraction of A is XA then, mole fraction of B can be expressed as (1-XA)

300XA=125(1-XA)

300XA=125- 125XA

300XA+125XA=125

425XA=125

XA=5/17

XB= (1-XA)=(1-5/17)

mole fraction of B= 12/17